1.State four of the transformations of energy which occurs at a power station which uses coal as its fuel. Account for energy losses.
2. A ball of mass1kg is dropped from a height of 7m and rebounds to a height of 4.5m. Calculate: (a) its Kinetic Energy just before impact.
(b) its initial rebound velocity and Kinetic Energy. Account for the loss of Kinetic Energy on impact.
3. State the energy transformation which occurs when a boy shoots a stone from a catapult to hit a target.
1
Expert's answer
2020-08-24T11:57:43-0400

1) The furnace heat converts boiler water to steam, which is then used to spin turbines that turn generators. Thus chemical energy stored in coal is converted successively into thermal energy, mechanical energy and, finally, electrical energy.

2)

a)


"K_1=mgH=(1)(9.8)(7)=69\ J"

b)


"K_2=mgh=(1)(9.8)(4.5)=44\ J"

"v_2=\sqrt{2gh}=\sqrt{2(9.8)(4.5)}=9.4\frac{m}{s}"

3)When the stone is released, the potential energy is transformed suddenly as the kinetic energy of motion of the projectile. When the projectile hits a target, the kinetic energy is used to do the work of crumbling a target.

Eventually all of this energy is dissipated as heat energy.